### v1

# On the Fourier Coefficients of Critical Gaussian Multiplicative Chaos

Abstract.
We continue the study of the Fourier coefficients of Gaussian multiplicative chaos (GMC) recently initiated by Garban and Vargas [16]. We show that the Fourier coefficients $c_{n}$ of critical GMC on the unit interval converge to zero in probability as the frequency $n$ tends to infinity, by establishing tightness for $\{(\log n)|c_{n}|\}_{n\geq 1}$.
1. Introduction
The theory of Gaussian multiplicative chaos (GMC) was initiated by Kahane [20] in the 1980s and gives meaning to measures formally written as
(1.1) $\mu_{\gamma}(\mathrm{d}x)\,``\hskip-2.00749pt=\hskip-2.00749pt"\,e^{\gamma X(x)-\tfrac{\gamma^{2}}{2}\mathbb{E}X(x)^{2}}\mathrm{d}x.$
Here, $X$ and is a logarithmically correlated Gaussian field on $\mathbb{R}^{d}$, $\gamma>0$ is an inverse temperature parameter, and $\mathrm{d}x$ denotes Lebesgue measure on some Borel set $D\subseteq\mathbb{R}^{d}$. Initially motivated by the need for a rigorous model for energy dissipation and intermittency in the Kolmogorov-Obukhov-Mandelbrot theory of developed turbulence [30, 22, 33], this theory has since found applications in many other fields, including finance [12], quantum gravity [14], the extreme values of log-correlated fields [8, 11, 6, 7] (see also [36] and the references therein) and related problems in random matrix theory and number theory (e.g., [41, 32, 17, 18, 38]).
The difficulty in defining (1.1) lies in the fact that such a field $X$ has a covariance kernel which is singular along the diagonal, being of the form
$K(x,y)=\mathbb{E}[X(x)X(y)]=\log_{+}\frac{1}{|x-y|}+f(x,y),\quad\log_{+}(x)=\max(x,0)$
for some smooth function $f:D\times D\longrightarrow\mathbb{R}$. It therefore cannot be defined point-wise and is instead a distribution (in the sense of Schwartz) with probability one. To rigorously define its exponential, one typically considers some regularisation $X_{\epsilon}$ of $X$ to eliminate the logarithmic singularity and defines $\mu_{\gamma}$ as the weak limit of
(1.2) $\mu_{\gamma,\epsilon}(\mathrm{d}x)=e^{\gamma X_{\epsilon}(x)-\frac{\gamma^{2}}{2}\mathbb{E}X_{\epsilon}(x)^{2}}\mathrm{d}x$
as $\epsilon\to 0$, which is known to be non-degenerate if $\gamma<\gamma_{c}=\sqrt{2d}$ [37, 5, 39]. At $\gamma=\gamma_{c}$, a non-trivial limit can still be obtained by considering the so-called derivative martingale [13]
(1.3) $\mu_{\gamma_{c},\epsilon}(\mathrm{d}x)=\big(\hskip-2.00749pt-\hskip-2.00749ptX_{\epsilon}(x)+\sqrt{2}\log(1/\epsilon)\big)e^{\gamma X_{\epsilon}(x)-\frac{\gamma^{2}}{2}\mathbb{E}X_{\epsilon}(x)^{2}}\mathrm{d}x$
or by multiplying (1.2) by a divergent factor of $\sqrt{\log(1/\epsilon)}$ [34, 19]. Both of these procedures yield the same limit up to a multiplicative constant, namely critical GMC, which is non-atomic and supported on a set of Hausdorff measure zero [35]. In this paper, our primary focus will be to study the asymptotic behavior of the Fourier coefficients of this measure.
1.1. Harmonic analysis of multifractal measures
Given a measure $\mu$ on the unit circle identified with $0,2\pi)$, one can ask whether or not its Fourier coefficients
$\hat{\mu}(n)=\int_{0}^{2\pi}e^{-2\pi in\theta}\mathrm{d}\mu(\theta)$
vanish at infinity. Measures satisfying this condition are known as Rajchman measures and form a class lying strictly between those of absolutely continuous and continuous measures (on $0,2\pi)$). That these inclusions are proper is a consequence of Menšov’s [31] famous construction of a singular Rajchman measure, and Wiener’s characterization of continuous measures [42] as those for which the $\hat{\mu}(n)$ vanish in mean, respectively. Rajchman measures can alternatively be characterized by their common null sets (meaning sets to which they assign zero measure)—which are an intermediate between countable sets and sets of Lebesgue measure zero (see [27, 26])—and are more generally related to the notion of sets of multiplicity and uniqueness in harmonic analysis [21].
In a recent work [16], Garban and Vargas initiated the study of the Fourier coefficients of (one-dimensional) GMC, proving in particular that the latter is Rajchman if $\gamma<\sqrt{2}$  [^1] To be precise, they prove this for the GMC arising from the Gaussian free field on the unit circle.. This resolved a longstanding open problem, and together with [15] was the first result of this kind for random multifractal measures. They also supplemented their result by a series of open problems that have since garnered some attention.
For instance, define the Fourier dimension of a measure $\mathrm{dim}_{F}(\mu)$ to be the largest $s$ for which $\hat{\mu}(n)\leq C/n^{s/2}$ for some constant $C>0$, and let $\mu_{\gamma}$ denote GMC on the unit circle with parameter $\gamma$. Using a fourth moment calculation, Garban and Vargas proved that $\mathrm{dim}_{F}(\mu)\geq(1-\gamma^{2}/2)/2$ if $\gamma<1/\sqrt{2}$, and conjectured that
(1.6) $\mathrm{dim}_{F}(\mu_{\gamma})=\left\{\begin{array}[]{ll}1-\gamma^{2}&\mbox{if }0<\gamma<1/\sqrt{2}\\ (\sqrt{2}-\gamma)^{2}&\mbox{if }1/\sqrt{2}\leq\gamma<\sqrt{2}.\end{array}\right$.
This would imply that $\dim_{F}(\mu_{\gamma})$ coincides with the correlation dimension of $\mu_{\gamma}$, defined as the maximal $s$ for which
$\iint_{[0,1]^{2}}\frac{\mu_{\gamma}(\mathrm{d}x)\mu_{\gamma}(\mathrm{d}y)}{|x-y|^{s}}<\infty\quad\text{a.s.},$
thereby making $\mu_{\gamma}$ a Salem measure.
This conjecture (1.6) was recently settled by Lin, Qiu and Tan [24, 25] in a series of works that develop a general framework to study the Fourier dimension of multiplicative chaos measures (including but not limited to one-dimensional GMC). The latter builds on a previous work of Chen, Han, Qiu and Wang [10], which used the theory of vector-valued martingales to compute the Fourier dimension of the Mandelbrot canonical cascade measure, answering a question of Mandelbrot from 1976.
By contrast, the picture for critical GMC is far murkier. It is well known the Fourier dimension of a measure is bounded above by its Hausdorff dimension, which implies that $\mathrm{dim}_{F}(\mu_{\gamma_{c}})=0$ when $\mu_{\gamma_{c}}$ is a critical GMC measure (i.e.  the in-probability limit of (1.3) as $\epsilon\to 0$). To the best of our knowledge, there are no further results in this direction, and the main open problem in [16] was to determine whether or not critical GMC is Rajchman.
Our goal is to provide a first step towards the resolution of this conjecture by proving the following.
1.2. Main results
In what follows, let $\mu=\mu_{\sqrt{2}}$ be the critical Gaussian multiplicative chaos measure arising from a $\star$–scale invariant field $X$ on $[0,1]$ (defined in Section 2), meaning
$\mu(\mathrm{d}\theta):=\lim_{t\to\infty}\mu_{t}(\mathrm{d}\theta),\quad\text{ where }\mu_{t}(\mathrm{d}\theta):=\sqrt{t}e^{\sqrt{2}X_{t}(\theta)-\mathbb{E}X_{t}(\theta)^{2}}\mathrm{d}\theta,$
where the limit is in probability, in the space of Radon measures on $[0,1]$. Let $c_{n}$ denote the $n$-th Fourier coefficient of $\mu$. Our main result establishes tightness for $\{(\log n)|c_{n}|\}_{n\geq 1}$, which implies that $c_{n}\to 0$ in probability as $n\to\infty$.
Theorem 1.1.
For any $\varepsilon>0$, there exists a $K=K(\varepsilon)>0$ such that
$\sup_{n\geq 1}\mathbb{P}\big\{(\log n)|c_{n}|>K\big\}<\varepsilon.$
The $(\log n)^{-1}$ scaling stems from what can loosely be described as a freezing phenomenon, which takes effect as soon as $\gamma\geq 1/\sqrt{2}$ and contributes logarithmic terms in $n$. This is explained in Section 1.3, which also contains a heuristic explanation of why we believe our bound to be sharp—that is, $|c_{n}|$ is indeed of the order of $1/\log n$.
Our proof is based on a second moment estimate for $c_{n}$ on a good event. While the approach is similar in spirit to Berestycki [5] and Lacoin’s [23] constructions of subcritical and critical GMC, respectively, the presence of an oscillating factor in the integral defining $c_{n}$ here introduces significant complications. We also require a more delicate notion of “good” points than in [5, 23], that uses insights from the study of the maximum of log-correlated Gaussian fields, following the works of Bramson–Ding–Zeitouni [7] and Madaule [28]. An overview of the proof is given in Section 1.3.
By comparison, Garban and Vargas bypass the explicit study of $c_{n}$ (when $\gamma>1/\sqrt{2}$) by appealing to the Riemann-Lebesgue lemma. To show that $c_{n}\to 0$ almost surely, they note that it suffices by the latter to show that $\mu_{\gamma}^{*d}$ is in $L^{1}$ for sufficiently large $d$ (where ∗d denotes $d$-fold convolution). Unfortunately, this approach breaks down at criticality as their argument would force one to take $d=\infty$.
1.3. Proof elements
Let $X$ be a $\star$–scale invariant field as above, and consider the multiplicative chaos measure $\mu_{\gamma}:=\lim_{t\to\infty}\mu_{t,\gamma}$ associated to $X$ for $\gamma\in(1/\sqrt{2},\sqrt{2}]$, where
(1.7) $\mu_{t,\gamma}(\mathrm{d}\theta):=\big(\sqrt{t}\cdot\mathbf{1}(\gamma=\sqrt{2})+\mathbf{1}(\gamma\neq\sqrt{2})\big)e^{\gamma X_{t}(\theta)-(\gamma^{2}/2)t}.$
Our goal is to estimate $\mathbb{E}[|c_{n,t}|^{2}]$, where $c_{n,t}:=\int e^{in\theta}{\mu_{\sqrt{2},t}(\mathrm{d}\theta)}.$
Before this, it is helpful to understand the second moment of the total mass $\int\mu_{\gamma,t}(\mathrm{d}\theta)$ for $\gamma<\sqrt{2}$, without the oscillating factor $e^{in\theta}$. Since $X_{t}$ is logarithmically correlated, Fubini’s theorem implies that
$\mathbb{E}\bigg[\bigg(\int\mu_{\gamma,t}(\mathrm{d}\theta)\bigg)^{2}\bigg]\approx\iint\frac{\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}}{|\theta_{1}-\theta_{2}|^{\gamma^{2}}},$
which is finite if and only if $\gamma^{2}<1$ (the so-called $L^{2}$ phase” of GMC). For larger $\gamma$, the expectation is inflated by points $\theta$ where $X_{t}(\theta)$ is atypically large at many scales $t$. To circumvent this, one usually restricts the integral to “good” points $\theta\in\mathcal{G}$ for which $(X_{t}(\theta),t\geq 0)$ is well-behaved [^2] This is the main idea in [5, 23], for instance. At $\gamma=\sqrt{2}$, this is achieved by working on the event that $\{\forall\theta\in[0,1]:\theta\in\mathcal{G}\}$—on which the bound
(1.8) $\iint\mu_{\sqrt{2},t}(\mathrm{d}\theta_{1})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{2})\leq\iint\mathbf{1}(\theta_{1},\theta_{2}\in\mathcal{G})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{1})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{2})$
holds by monotonicity, and the right-hand side has finite expectation.
An evident hurdle that arises when trying to adapt this approach to
(1.9) $|c_{n,t}|^{2}=\iint e^{in(\theta_{2}-\theta_{1})}\mu_{\sqrt{2},t}(\mathrm{d}\theta_{1})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{2})$
is that one cannot directly bound as in (1.8), due to the factor $e^{in(\theta_{2}-\theta_{1})}$. In particular, the set $\mathcal{G}$ must be picked so as to guarantee the positivity and real-valuedness of
(1.10) $\iint e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\big(\theta_{1},\theta_{2}\in\mathcal{G}\big)\mu_{\sqrt{2},t}(\mathrm{d}\theta_{1})\mu_{\sqrt{2},t}(\mathrm{d}\theta_{2}).$
Furthermore, if we hope to get savings from the oscillations in the integral above, we need some control over the regularity of
$\mathbb{E}\big[e^{\sqrt{2}(X_{t}(\theta_{1})+X_{t}(\theta_{2}))}\mathbf{1}(\theta_{1},\theta_{2}\in\mathcal{G})\big]$
as a function of the gap $\Delta=(\theta_{2}-\theta_{1})$. This further limits the choice of $\mathcal{G}$ to one for which the event $\{\theta_{1},\theta_{2}\in\mathcal{G}\}$ is relatively simple, while being restrictive enough to make the expectation of (1.10) of the correct order (at the very least, compensating for the additional $\sqrt{t}$ normalization present at criticality (1.7)).
We now explain where the $(\log n)^{-1}$ in Theorem 1.1 comes from, and why we believe it to be sharp. To begin with, a variant of the argument above reveals that the dominant contribution to (1.9) comes from $\log|\Delta|^{-1}\asymp\log n$. For concreteness, consider the contribution from $\Delta\in[n^{-100},n^{-1}]$, which is
(1.11) $\asymp te^{-2t}\iint_{|\Delta|\in[n^{-100},n^{-1}]}e^{\sqrt{2}(X_{t}(\theta_{1})+X_{t}(\theta_{2}))}\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}.$
(We ignore the $e^{in\Delta}$ factor since $n\Delta\leq 1$ here.)
Since $X_{t}$ is logarithmically correlated, one can show that for any $\theta_{1},\theta_{2}$
$X_{s}(\theta_{1})\approx X_{s}(\theta_{2}),\quad\forall s\leq r_{\Delta}:=\log|\Delta|^{-1},$
with high probability (see, e.g., Lemma 2.1), and it follows that (1.11) is
$\approx te^{-2t}\iint e^{2\sqrt{2}X_{r_{\Delta}}(\theta_{1})}e^{\sqrt{2}(X_{t}(\theta_{1})-X_{r_{\Delta}}(\theta_{1}))}e^{\sqrt{2}(X_{t}(\theta_{2})-X_{r_{\Delta}}(\theta_{2}))}\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}.$
Furthermore, $e^{\sqrt{2}(X_{t}(\theta_{i})-X_{r_{\Delta}}(\theta_{i}))}$ will typically be of the order of $e^{t-r_{\Delta}}/\sqrt{t-r_{\Delta}}$, which will allow us to bound the integral above by
(1.12) $\int_{n^{-100}}^{n^{-1}}e^{-2r_{\Delta}}\bigg(\int_{0}^{1}e^{2\sqrt{2}X_{r_{\Delta}}(\theta_{1})}\mathrm{d}\theta_{1}\bigg)\mathrm{d}\Delta$
as $t\to\infty$ (noting the change of variables and that $r_{\Delta}$ is bounded in $t$).
For large $n$, we recognize the integral over $\theta_{1}$ as nothing but the (approximate, unnormalized) total mass of a supercritical chaos measure, this time of parameter $\tilde{\gamma}=2\sqrt{2}$. [^3] The emergence of such a chaos was also found by Garban and Vargas (see Theorem 1.3 in [16]), where it is subcritical since they only consider $2\gamma<\sqrt{2}$. Recalling that $r_{\Delta}=\log|\Delta|^{-1}$, this is known to have order
$(\log|\Delta|^{-1})^{-(3\tilde{\gamma}/(2\sqrt{2}))}e^{r_{\Delta}(\tilde{\gamma}^{2}/2-(\tilde{\gamma}/\sqrt{2}-1)^{2})}=e^{3r_{\Delta}}(\log|\Delta|^{-1})^{-3}$
(see Theorem 2.2 in [29]). Inserting this into (1.12), we conclude that when $t\to\infty$,
$|c_{n,t}|^{2}\leq C\int_{n^{-100}}^{n^{-1}}\frac{1}{(\log|\Delta|^{-1})^{3}}\frac{\mathrm{d\Delta}}{\Delta}=C\int_{\log n}^{100\log n}\frac{\mathrm{d}u}{u^{3}}=O\big((\log n)^{-2}\big)$
with high probability, for some constant $C>0$. In particular, this suggests a non-trivial contribution from scales $\Delta\approx n^{-a}$, even when $a>0$ is large (see Remark 4.4).
1.4. Conjectures for subcritical chaos
While we do not attempt to make the heuristics from the previous section rigorous for $\gamma\in(1/\sqrt{2},\sqrt{2})$, they can nonetheless be used to make predictions in that range. There, the analogue of the integral in (1.12) contains the total mass of a chaos with parameter $\tilde{\gamma}=2\gamma$ and is therefore of the order of
$\int_{n^{-100}}^{n^{-1}}e^{-\gamma^{2}r_{\Delta}}\bigg(\frac{e^{r_{\Delta}(2\gamma^{2}-(\sqrt{2}\gamma-1)^{2})}}{(\log|\Delta|^{-1})^{3\gamma/\sqrt{2}}}\bigg)\mathrm{d}\Delta.$
By contrast with the critical case, the mass in this integral is concentrated around $\Delta\approx n^{-1}$, and we can factor out the correct normalization via the substitution $u=n\Delta$. What remains is approximately the integral of $1/|u|^{\kappa}$ over $[0,1]$ for some $\kappa<1$, which is integrable at zero.
Conjecture 1.2.
Let $\gamma\in(1/\sqrt{2},\sqrt{2})$, $\mu_{\gamma}$ denote the Gaussian multiplicative chaos measure on $[0,1]$ with parameter $\gamma$, and $c_{n}$ its $n$-th Fourier coefficient. Then
$(\log n)^{3\gamma/(2\sqrt{2})}n^{(\sqrt{2}-\gamma)^{2}/2}|c_{n}|$
converges in distribution to a (non-trivial) limiting random variable as $n\to\infty$.
Remark 1.3.
We recognize the Fourier dimension of $\mu_{\gamma}$ in the exponent of $n$ here.
Lastly, when $\gamma=1/\sqrt{2}$, the chaos of parameter $2\gamma=\sqrt{2}$ is critical and the heuristic above yields the following prediction.
Conjecture 1.4.
Let $\mu_{\gamma}$ denote the Gaussian multiplicative chaos measure on $[0,1]$ with parameter $\gamma=1/\sqrt{2}$, and $c_{n}$ its $n$-th Fourier coefficient. Then
$(\log n)^{1/4}n^{1/4}|c_{n}|$
converges in distribution to a (non-trivial) limiting random variable as $n\to\infty$.
Remark 1.5.
As explained in [16], Section 1.2.3, $n^{1/4}c_{n}$ is expected to behave like the $n$-th Fourier coefficient of the Holomorphic Multiplicative Chaos (HMC), which is known to be of order $(\log n)^{-1/4}$ [40].
1.5. Organisation and notation
The paper is organised as follows. Section 2 contains various preliminary results on logarithmically correlated fields needed for the proof of Theorem 1.1, which is then proved in Sections 3 and 4.
We will use standard asymptotic notation, writing $f(T)=o(g(T))$ to mean that $|f(T)/g(T)|\to_{T\to\infty}0$ and $f(T)=O(g(T))$ or $f(T)\ll g(T)$ to mean that $\limsup_{T\to\infty}|f(T)/g(T)|$ is bounded. A subscripted parameter next to $o$, $O$ or $\ll$ means that the implicit constant is allowed to depend on said parameter.
We also use
$(a\land b):=\min(a,b),\quad(a\lor b)=\max(a,b),$
and adopt the convention $\log x=\infty$ when $x<0$ so that, for instance, $\log(x)\land 1=1$ in this case.
2. Preliminary estimates
2.1. Properties of $\star$–scale invariant fields.
We will work with chaos measures associated to a $\star$–scale invariant field $X$ on $[0,1]$ (viewed as being 11–periodic) with $\star$–scale cut–off approximations $(X_{t};t\geq 0)$ (see [35]). Our setup is essentially the same as the one in [13] for $d=1$, and we refer to [3] for the construction of an approximating sequence $X_{t}$ satisfying the properties below.
In what follows, $(X_{t})_{t\geq 0}$ will denote a family of coupled, centered Gaussian fields on $[0,1]$ with covariances given by
(2.1) $\forall\theta_{1}<\theta_{2},\quad\mathbb{E}[X_{t}(\theta_{1})X_{s}(\theta_{2})]=K_{s\land t}(\theta_{2}-\theta_{1})=\int_{1}^{e^{s\land t}}\frac{k(u(\theta_{2}-\theta_{1}))}{u}\mathrm{d}u,$
for some fixed kernel $k\in C^{1}(\mathbb{R})$ associated to $X$, where $s\land t=\min(s,t)$. For convenience, we will assume that $k(0)=1$, $k$ is even and that it vanishes outside of the unit ball around the origin. The following properties then hold by construction:
  1. (1)
    $\mathbb{E}[X_{t}(\theta)^{2}]=t$ for every $\theta\in[0,1]$.
  2. (2)
    For every $\theta\in[0,1]$, $(X_{t}(\theta);t\geq 0)$ has independent increments, making $t\mapsto X_{t}(\theta)$ a standard Brownian motion $B_{t}$ starting from zero for every $\theta$.
  3. (3)
    $(X_{t}(\theta_{1})-X_{t_{0}}(\theta_{1});t\geq t_{0})$ is independent from $(X_{t}(\theta_{2})-X_{t_{0}}(\theta_{2});t\geq t_{0})$ for any $\theta_{1},\theta_{2}$ satisfying $|\theta_{1}-\theta_{2}|\geq e^{-t_{0}}$,
In particular, we have the following description for the joint law of $X_{t}$ at two different points.
Lemma 2.1 (Lemma 16 in [13]).
For any $\theta_{1}<\theta_{2}$ (both in $[0,1]$) and any $t_{0}\geq 0$, the law of $(X_{s}(\theta_{1}))_{s\leq t_{0}}$ admits the decomposition,
$X_{s}(\theta_{1})=P_{s}^{\theta_{2},\theta_{1}}+Z_{s}^{\theta_{2},\theta_{1}}$
where
$P_{s}^{\theta_{2},\theta_{1}}=-\int_{0}^{s}K_{u}^{\prime\prime}(\theta_{2}-\theta_{1})X_{u}(\theta_{2})\mathrm{d}u+K_{s}^{\prime}(\theta_{2}-\theta_{1})X_{s}(\theta_{2})$
is measurable with respect to $\sigma((X_{s}(\theta_{2})_{s\leq t_{0}})$, and $(Z_{s}^{\theta_{2},\theta_{1}})_{0\leq s\leq t_{0}}$ is a centered Gaussian field independent from the latter, with covariance kernel
$\mathbb{E}\big[Z_{s^{\prime}}^{\theta_{2},\theta_{1}}Z_{s}^{\theta_{2},\theta_{1}}\big]=s\land s^{\prime}-\int_{0}^{s\land s^{\prime}}K_{u}^{\prime}(\theta_{2}-\theta_{1})^{2}\mathrm{d}u.$
This essentially means that for any $\theta_{1},\theta_{2}$, $X_{k}(\theta_{1})$ and $X_{k}(\theta_{2})$ are almost perfectly correlated up to time $k=\log 1/|\theta_{1}-\theta_{2}|$ (and evolve independently after that time, as per property (3) above). For our purposes, we will only need the following estimates reflecting this correlation structure.
Lemma 2.2.
Let $k^{\prime}||_{\infty}=\sup_{x\in[-1,1]}|k^{\prime}(x)|$. Fix $t>0$, $r\in[0,t]$ and $s\in[0,t-r]$. Then the following estimates hold uniformly in $\Delta\in[0,e^{-r+1}]$,
[…]
Proof.
Assume for now that $\Delta\leq e^{-r}$. Then by expanding $k$,
[…]
For the second claim, we make the change of variables $v=e^{-r}u$ to write
$K_{s+r}(\Delta)-K_{r}(\Delta)=\int_{1}^{e^{s}}\frac{k(e^{r}\Delta v)}{v}\mathrm{d}v=\int_{1}^{e^{s}\land(\Delta e^{r})^{-1}}\frac{k(e^{r}\Delta v)}{v}\mathrm{d}v,$
where in the second equality, we used the fact that $k$ is supported on [1,1][-1,1]. Expanding $k$ as in Equation (2.6) yields
[…]
and (2.5) then follows since $\Delta\leq e^{-r}$. The claim for $\Delta\geq e^{-r}$ follows from Equations (2.2) and (2.5) for $\Delta=e^{-r}$, and the estimate
[…]
For any times $t>r>0$, we adopt the notation
(2.7) $X_{t}^{(r)}:=X_{t+r}-X_{r},$
noting that the field $(X_{t}^{(r)}(\theta))_{\theta\in[0,1]}$ is $\star$-scale invariant with kernel $k^{(r)}(x):=k(e^{r}x)$ since
$\mathbb{E}[X_{s}^{(r)}(\theta_{1})X_{s}^{(r)}(\theta_{2})]=\bigg(\int_{1}^{e^{s+r}}-\int_{1}^{e^{s}}\bigg)\frac{k(u(x-y))}{u}\mathrm{d}u=\int_{1}^{e^{s}}\frac{k^{(r)}(u(x-y))}{u}\mathrm{d}u.$
Our argument will require some control on the large values of $X_{r_{n}}$ and $(X_{t}^{(r_{n})})_{t\geq 0}$, where $r_{n}$ is a carefully chosen time depending on the frequency $n$ of the Fourier coefficient being studied (see Proposition 4.3). Following [13, 28], the typical size of $\max_{\theta\in[0,1]}X_{t}(\theta)$ fluctuates around
(2.8) $m(t):=\sqrt{2}t-\frac{3}{2\sqrt{2}}\log t$
as made precise by the following proposition.
Proposition 2.3 (Theorem 1.1 in [1]).
Let $m(t)$ be as in Equation (2.8). Then the family
$\Big(\max_{\theta\in[0,1]}X_{t}(\theta)-m(t)\Big)_{t\geq 1},$
is tight.
This is what one expects from any log–correlated field (see, e.g. [4]). In fact, one can typically go further and show that the paths $s\mapsto X_{s}(\theta),s\leq t$ of points $\theta$ for which $X_{t}(\theta)$ is near the maximum cannot fluctuate too wildly. Instead, they display almost linear growth while remaining under a suitably chosen barrier function. This phenomenon was first leveraged by Bramson [9] to study the maximal particle of branching Brownian motion, and underpins most of the works that followed on the extrema of other log–correlated fields [7, 8, 2, 11].
In what follows, we will use the following upper bound essentially due to Lacoin [23].
Proposition 2.4 (Proposition 2.4 in [23]).
Let $(X_{t})_{t\geq 0}$ be defined as in Equation (2.1). Then almost surely,
$\sup_{t\geq e}\sup_{\theta\in[0,1]}\bigg(X_{t}(\theta)-\sqrt{2}t+\frac{\log t}{2\sqrt{2}}-\frac{4\log\log t}{\sqrt{2}}\bigg)<\infty.$
It follows that with high probability, the fields $(X_{t}(\theta))_{t\geq 0}$ remain below
(2.9) $U(t):=\begin{cases}\sqrt{2}t-\frac{\log t}{2\sqrt{2}}+\frac{4\log\log t}{\sqrt{2}}&\quad\text{if }s\geq e,\\ \infty&\quad\text{if }s<e.\end{cases}$
at every point $\theta\in[0,1]$ in the following sense.
Corollary 2.5.
Let $U(t)$ be defined as in Equation (2.9). Then
$\mathbb{P}\Big(\forall\theta\in[0,1],\forall t\geq 0,X_{t}(\theta)\leq U(t)+A\Big)\underset{\text{$A\to\infty$}}{\longrightarrow}1.$
2.2. Brownian estimates
We record the following estimates for the probability that a Brownian motion stays under a barrier starting at time $e$ (in light of (2.9)). Such events will arise from restricting to samples $\omega$ for which the event in Corollary 2.5 holds.
Lemma 2.6.
Let $B_{t}$ be a standard linear Brownian motion and $a>0$. Then for any $t>e$,
(2.10) $\mathbb{P}\Big(\sup_{s\in[e,t]}B_{s}\leq a\Big)\ll\frac{(a+1)}{\sqrt{t}+1}.$
Furthermore, if $a,b>0$ and $\mathbb{P}_{(0,a)}^{(t,b)}$ denotes the law of a Brownian bridge of length $t>e$ from $a$ to $b$, then
(2.11) $\mathbb{P}_{(0,a)}^{(t,b)}\Big(B_{s}\geq 0,\forall s\in[e,t]\Big)\ll\frac{(a+1)(b+1)}{t+1}.$
Proof.
Both inequalities are a direct consequence of the reflection principle after integrating over the value of $B_{e}$. ∎
3. Proof of the main theorem
Let $X$ be a $\star$-scale invariant field with approximations $(X_{t})_{t\geq 0}$, and define its associated (critical) GMC measure $\mu$ as
$\mu(\mathrm{d}\theta):=\lim_{t\to\infty}\mu_{t}(\mathrm{d}\theta),\quad\text{ where }\mu_{t}(\mathrm{d}\theta):=\sqrt{t}e^{\sqrt{2}X_{t}(\theta)-\mathbb{E}X_{t}(\theta)^{2}}\mathrm{d}\theta.$
This limit holds in probability, and it follows that
[…]
We want to show that $\{(\log n)\cdot|c_{n}|\}_{n\geq 1}$ is tight, which we do by proving an upper bound of the correct order for the expectation of
(3.2) $|c_{n,t}|^{2}=\iint_{[0,1]^{2}}e^{in(\theta_{2}-\theta_{1})}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2}).$
on a good event $\mathcal{E}_{n,t}^{(A)}$. This is the content of Proposition 3.1 below. As a byproduct of its proof, we identify the set of pairs $(\theta_{1},\theta_{2})\in[0,1]^{2}$ over which the integral in (3.2) yields a leading order contribution, and those for which the contribution is negligible in the limit (see Remark 4.4).
Proposition 3.1.
Let $A>0$ and $0<\delta<1/4$. Let $r_{n}:=\delta\log n$. For any $t\geq r_{n}$, consider the event
$\mathcal{E}^{(A)}_{n,t}:=\big\{\forall\theta\in[0,1],\,\theta\in\mathcal{G}_{n,t}^{(A)}\big\},$
where
(3.3) $\mathcal{G}_{n,t}^{(A)}=\left\{\theta\in[0,1]:\,\,\begin{aligned} &X_{s}^{(r_{n})}(\theta)\leq U(s)+A,\,\forall s\in[0,t-r_{n}]\\ &X_{r_{n}}(\theta)\leq m(r_{n})+A\end{aligned}\right\}$
and $m(t),U(t)$ are defined in Equations (2.8), (2.9) respectively. Then there exists an integer $N>2$ such that for any $n\geq N$,
$\limsup_{t\to\infty}\mathbb{E}\Big[|c_{n,t}|^{2};{\mathcal{E}_{n,t}^{(A)}}\Big]\ll_{A}(\log n)^{-2},$
where the implicit constant is uniform in $n$.
The choice of good event $\mathcal{E}_{n,t}^{(A)}$ will be explained in Section 4. For the time being, we show how Theorem 1.1 follows straightforwardly from this proposition.
Proof of Theorem 1.1 assuming Proposition 3.1.
Noting that $|c_{n}|\leq\mu([0,1])$ is almost surely finite, it suffices to prove that $\{|c_{n}|\log n\}_{n\geq N}$ is tight where $N$ is the constant from Proposition 3.1.
Fix $\epsilon>0$ and let $K>0$ be arbitrary for now. Then by (3.1) and the Portmanteau theorem, we have that
$\mathbb{P}\Big(|c_{n}|(\log n)>K\Big)$ $\leq\limsup_{t\to\infty}\mathbb{P}\Big(|c_{n,t}|(\log n)>K\Big)$
for any $n>0$. Furthermore, by definition of $\mathcal{E}_{n,t}^{(A)}$, the pre-limit quantity on the right-hand side is bounded above by
$\mathbb{P}\Big(\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}$ $|c_{n,t}|(\log n)>K\Big)+\mathbb{P}\Big(\exists\theta\in[0,1]:X_{r_{n}}(\theta)>m(r_{n})+A\Big)$
(3.4) $+\mathbb{P}\Big(\exists s\in[0,\infty),\exists\theta\in[0,1]:X_{s}^{(r_{n})}(\theta)>U(s)+A\Big),$
for any $A>0$. We can then make the second and third terms in (3) arbitrarily small in a way that is uniform in $n$, by taking $A$ large enough. Indeed, for any $n$, $(X_{t}^{(r_{n})})_{t\geq 0}$ (defined in (2.7)) is a sequence of cutoff approximations to a $\star$–scale invariant field with variance $t$ at every point. Corollary 2.5 and Proposition 2.3 thus imply that
(3.5) $\sup_{n>2}\mathbb{P}\Big(\exists s\in[0,\infty),\exists\theta\in[0,1]:X_{s}^{(r_{n})}(\theta)>U(s)+A\Big)<\varepsilon/3,$
(3.6) $\sup_{n>2}\mathbb{P}\Big(\exists\theta\in[0,1]:X_{r_{n}}(\theta)>m(r_{n})+A\Big)<\varepsilon/3,$
for sufficiently large $A$. For such an $A$, it follows that
$\sup_{n\geq N}\mathbb{P}\Big(|c_{n}|(\log n)>K\Big)$ $\leq\sup_{n\geq N}\limsup_{t\to\infty}\mathbb{P}\Big(\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}|c_{n,t}|(\log n)>K\Big)+2\varepsilon/3.$
Using Chebyshev’s inequality and Proposition 3.1, we can then write
$\sup_{n\geq N}\limsup_{t\to\infty}\mathbb{P}\Big(\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}|c_{n,t}|(\log n)>K\Big)$
$\leq\sup_{n\geq N}\frac{(\log n)^{2}}{K^{2}}\limsup_{t\to\infty}{\mathbb{E}\big[|c_{n,t}|^{2};\mathcal{E}_{n,t}^{(A)}\big]}\ll_{A}\frac{1}{K^{2}}.$
Tightness follows by picking $K=K(A)$ to make the left-hand side $\varepsilon/3$. ∎
What’s left is to prove Proposition 3.1, which is the main difficulty here.
4. Proof of Proposition 3.1
4.1. Initial reductions
To bound $|c_{n,t}|\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}$, we first want to extract a local condition (meaning a condition on each pair of points $\theta_{1},\theta_{2}$) in (3.2) from the global event $\mathcal{E}_{n,t}^{(A)}$, noting that one cannot directly upper bound the integrand due to the presence of the oscillating factor $e^{in\Delta}$. We instead note the following pointwise inequality on the sample space
(4.1) $\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}|c_{n,t}|^{2}$ $=\mathbf{1}_{\mathcal{E}_{n,t}^{(A)}}\iint_{[0,1]^{2}}e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\Big\{\theta_{1},\theta_{2}\in\mathcal{G}_{n,t}^{(A)}\Big\}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2})$
$\leq\iint_{[0,1]^{2}}e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\Big\{\theta_{1},\theta_{2}\in\mathcal{G}_{n,t}^{(A)}\Big\}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2}).$
We used the fact that the indicator in the integrand is redundant in the first line and positivity of the double integral on the second, noting that it equals the modulus squared of the nn–th Fourier coefficient of
$\tilde{\mu}_{t}(\mathrm{d}\theta):=\mathbf{1}\Big\{\theta\in\mathcal{G}_{n,t}^{(A)}\Big\}\mu_{t}(\mathrm{d}\theta).$
It therefore suffices to prove the analogous claim (to Proposition 3.1) for the Fourier coefficients of $\tilde{\mu}_{t}$.
To that end, fix $0<\delta<1/4$ and partition $0,1]^{2}$ into the following two ranges:
(4.2) $\mathrm{(I)}$ $=\{(\theta_{1},\theta_{2}):|\theta_{2}-\theta_{1}|\in[\Delta_{n},1]\}$
(4.3) $\mathrm{(II)}$ $=\{(\theta_{1},\theta_{2}):|\theta_{2}-\theta_{1}|\in[0,\Delta_{n})\},$
where
$\Delta_{n}=:en^{-\delta},\text{ noting that }\log\Delta_{n}^{-1}=r_{n}-1.$
Denote the contribution to (3.2) coming from each of these regions by
(4.4) $\mathcal{C}_{n,t}(\#):=\iint_{(\mathrm{\#})}e^{in(\theta_{2}-\theta_{1})}\mathbf{1}\Big\{\theta_{1},\theta_{2}\in\mathcal{G}_{n,t}^{(A)}\Big\}\mu_{t}(\mathrm{d}\theta_{1})\mu_{t}(\mathrm{d}\theta_{2})$
for $\in\{\mathrm{I},\mathrm{II}\}$.
The restriction on $\delta$ is technical, and is necessary for our bound for the contribution from (I) (see (4.17)) to be of the correct order. On (II), we will ignore the $e^{in(\theta_{2}-\theta_{1})}$ by the triangle inequality. While this may seem excessively inefficient—since $\delta<1$, $e^{in(\theta_{2}-\theta_{1})}$ will still oscillate with nn—it will become apparent that any $\delta>0$ yields the same bound (due to the observation in Remark 4.4).
4.2. The oscillatory range
We first study the contribution from $(\mathrm{I)}$ where $e^{in\Delta}$ oscillates very rapidly.
Proposition 4.1.
Let $A>0$. Then
(4.5) $\lim_{t\to\infty}\mathbb{E}\big[\mathcal{C}_{n,t}(\mathrm{I})\big]=o_{A}\big((\log n)^{-2}\big)$
for large enough $n$.
Proof.
Recall that $r_{n}=\delta\log n$ where $0<\delta<1/4$. By properties (2) and (3) in Section 2, we know that for any pair $\theta_{1},\theta_{2}\in\mathrm{(I)}$, the processes
(4.6) $X_{s}^{(r_{n})}(\theta_{1})\text{ and }X_{s}^{(r_{n})}(\theta_{2})\text{ for }{s\leq t-r_{n}}$
are independent, identically distributed, and independent from the $\sigma$–algebra generated by the field $X_{s}$ up to time $s\leq r_{n}$. Furthermore, the indicator $\mathbf{1}\big\{\theta_{1},\theta_{2}\in\mathcal{G}_{n,t}^{(A)}\big\}$ factors into
$\mathbf{1}\Big\{X_{r_{n}}(\theta_{i})\leq m(r_{n})+A,i\in\{1,2\}\Big\}$
$\times\mathbf{1}\Big\{X_{s}^{(r_{n})}(\theta_{i})\leq U(s)+A,\forall s\in[0,t-r_{n}],i\in\{1,2\}\Big\}$
$:=\mathcal{I}_{\leq r_{n}}\mathcal{I}_{>r_{n}}$
where the $\mathcal{I}_{\leq r_{n}}$ is measurable with respect to (the $\sigma$-algebra generated by) $X_{r_{n}}$, and $\mathcal{I}_{>r_{n}}$ with respect to $(X^{(r_{n})}_{s})_{s\in[0,t-r_{n}]}$.
It follows that for any $(\theta_{1},\theta_{2})$,
(4.7) $\mathbb{E}\Big[e^{\sqrt{2}(X_{t}(\theta_{1})+X_{t}(\theta_{2}))}\mathbf{1}\Big\{\theta_{1},\theta_{2}\in\mathcal{G}_{n,t}^{(A)}\Big\}\Big]$
$=\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))}\mathcal{I}_{\leq r_{n}}\Big]\mathbb{E}\bigg[\prod_{i=1}^{2}e^{\sqrt{2}(X_{t}(\theta_{i})-X_{r_{n}}(\theta_{i}))}\mathcal{I}_{>r_{n}}\bigg]$
and crucially, for $(\theta_{1},\theta_{2})\in(\mathrm{I})$ the rightmost factor is independent of $(\theta_{1},\theta_{2})$ since $\log|\theta_{2}-\theta_{1}|\leq r_{n}-1\leq r_{n}$. By translation invariance of the law of $(X_{s}^{(r_{n})})_{s\geq 0}$, the factor in question equals
$\mathbb{E}\Big[e^{\sqrt{2}X_{t-r_{n}}^{(r_{n})}(0)}\mathcal{I}_{>r_{n}}\Big]^{2}.$
Using $U(s)\leq\sqrt{2}s$,
$\mathcal{I}_{>r_{n}}\leq\mathbf{1}\big\{X_{s}^{(r_{n})}(0)\leq\sqrt{2}s+A,\forall s\in[0,t-r_{n}]\big\},$
and we conclude by Fubini’s theorem that
$\mathbb{E}[\mathcal{C}_{n,t}(\mathrm{I})]$ $\leq e^{-2r_{n}}\iint_{(\mathrm{I})}e^{in(\theta_{2}-\theta_{1})}\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))}\mathcal{I}_{\leq r_{n}}\Big]\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}$
(4.8) $\times{t}e^{-2t+2r_{n}}\mathbb{E}\Big[e^{\sqrt{2}X_{t-r_{n}}^{(r_{n})}(0)}\mathbf{1}\big\{X_{s}^{(r_{n})}(0)\leq\sqrt{2}t+A,\forall s\in[0,t-r_{n}]\big\}\Big]^{2}.$
We begin by showing that the factor in (4.2) is bounded in $t$ and $n$. Under the measure $\mathbb{Q}$ defined through
$\frac{\mathrm{d}\mathbb{Q}}{\mathrm{d}\mathbb{P}}={e^{\sqrt{2}X_{t-r_{n}}^{(r_{n})}(0)-\mathbb{E}[X_{t-r_{n}}^{(r_{n})}(0)^{2}]}},$
the mean of $(X_{s}^{(r_{n})}(0),s\in[0,t-r_{n}]$) is shifted by $\sqrt{2}s$ by Girsanov’s theorem. The quantity on the second line in (4.2) is thus bounded by
$\limsup_{t\to\infty}t\cdot\mathbb{P}\Big(X_{s}^{(r_{n})}(0)\leq A,\,\forall s\in[e,t-r_{n}]\Big)^{2}$
$\ll_{A}\limsup_{t\to\infty}\frac{t}{t-r_{n}}=O_{A}(1),$
by Lemma (2.10).
It remains to show that
$\limsup_{t\to\infty}e^{-2r_{n}}\iint_{(\mathrm{I})}e^{in(\theta_{2}-\theta_{1})}\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))}\mathcal{I}_{\leq r_{n}}\Big]\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}=o_{A}\big((\log n)^{-2}\big).$
For fixed $\theta_{1},\theta_{2}$, consider the change of measure given by
(4.9) $\frac{\mathrm{d}\mathbb{Q}_{\Delta}}{\mathrm{d}\mathbb{P}}=e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))-\mathbb{E}[(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))^{2}]}.$
The mean of $X_{r_{n}}(\theta_{1})$ under $\mathbb{Q}_{\Delta}$ is then shifted by
(4.10) $\sqrt{2}\cdot\text{Cov}\Big(X_{r_{n}}(\theta_{1}),X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2})\Big)=\sqrt{2}\big(r_{n}+K(\Delta)\big)$
where $\Delta=\theta_{2}-\theta_{1}$ and $K=K_{r_{n}}$ (cf.  Equation (2.1). The same is true for $X_{r_{n}}(\theta_{2})$ since $K(\Delta)=K(|\Delta|)$. Noting that
$\mathbb{E}[(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))^{2}]=2r_{n}+2K_{r_{n}}(|\theta_{2}-\theta_{1}|),$
the change of variables $(\theta,\Delta)=(\theta_{1},\theta_{2}-\theta_{1})$ thus yields
$e^{-2r_{n}}\iint_{(\mathrm{I})}e^{in(\theta_{2}-\theta_{1})}\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))}\mathcal{I}_{\leq r_{n}}\Big]\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}$
$=2\int_{\Delta_{n}}^{1}e^{in\Delta+2K(\Delta)}\int_{0}^{1}\mathbb{Q}_{\Delta}\mathbf{(}\mathcal{I}_{\leq r_{n}})\mathrm{d}\theta\mathrm{d}\Delta$
$=2\int_{\Delta_{n}}^{1}e^{in\Delta+2K(\Delta)}\mathbb{Q}_{\Delta}\mathbf{(}\mathcal{I}_{\leq r_{n}})\mathrm{d}\Delta$
where $\Delta_{n}=en^{-\delta}$.
We integrate by parts to exploit the cancellation coming from $e^{in\Delta}$. Doing so yields
$\int_{\Delta_{n}}^{1}$ $e^{in\Delta}e^{2K(\Delta)}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\mathrm{d}\Delta$
(4.11) $=\frac{e^{in\Delta}e^{2K(\Delta)}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})}{in}\bigg|^{1}_{\Delta_{n}}-\int_{\Delta_{n}}^{1}\frac{e^{in\Delta}\partial_{\Delta}\big(e^{2K(\Delta)}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\big)}{in}\mathrm{d}\Delta.$
Noting that $\mathbb{Q}_{\Delta}$ is a probability measure, the first term is
$\ll\frac{e^{2K(\Delta_{n})}+e^{2K(1)}}{n}\ll\frac{e^{2r_{n}}}{n}\leq n^{2\delta-1}=o\big((\log n)^{-2}\big).$
By triangle inequality, the second term in (4.11) is
$\ll\frac{1}{n}\sup_{\Delta\in[\Delta_{n},1]}\Big(|K^{\prime}(\Delta)|e^{2K(\Delta)}+e^{2K(\Delta)}\partial_{\Delta}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\Big)$
(4.12) $\ll\frac{e^{2K(\Delta_{n})}}{n}\bigg(\sup_{\Delta\in[\Delta_{n},1]}|K^{\prime}(\Delta)|+\sup_{\Delta\in[\Delta_{n},1]}\partial_{\Delta}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\bigg).$
Using the fact that $e^{2K(\Delta_{n})}\leq n^{2\delta}$ and that $K^{\prime}$ is uniformly bounded on $\Delta_{n},1]$ by
[…]
(cf.  Equation (2.1)) we conclude that (4.12) is
(4.14) $\ll n^{3\delta-1}+n^{2\delta-1}\sup_{\Delta\in[\Delta_{n},1]}\partial_{\Delta}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}}).$
Recalling the mean shift in Equation (4.10),
$\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})=\mathbb{P}\Big(X_{i}\leq-\sqrt{2}K(\Delta)-\frac{3}{2\sqrt{2}}\log r_{n}+A,\,\forall i\in\{1,2\}\Big)$
where $X_{1},X_{2}$ are centered Gaussian random variables with mean 0, variance $r_{n}$ and covariance $K(\Delta)$. Equivalently, by Fubini’s theorem, this equals
$\int_{-\infty}^{H(\Delta)}\int_{-\infty}^{H(\Delta)}p_{\Delta}(x_{1},x_{2})\mathrm{d}x_{1}\mathrm{d}x_{2}$
where $p_{\Delta}$ is the density of $(X_{1},X_{2})$ and $H(\Delta)=-\sqrt{2}K(\Delta)-\frac{3}{2\sqrt{2}}\log r_{n}+A$ for simplicity. We differentiate by repeated applications of Leibniz’s rule, whose validity follows from the uniform bounds established later in the proof. This yields
$H^{\prime}(\Delta)\bigg(\int_{-\infty}^{H(\Delta)}p_{\Delta}\big(x_{1},H(\Delta)\big)\mathrm{d}x_{1}+\int_{-\infty}^{H(\Delta)}p_{\Delta}\big(H(\Delta),x_{2}\big)\mathrm{d}x_{2}\bigg)$
(4.15) $+\iint_{(-\infty,H(\Delta)]^{2}}\partial_{\Delta}p_{\Delta}(x_{1},x_{2})\mathrm{d}x_{1}\mathrm{d}x_{2}.$
The boundary terms in (4.2) are equal by symmetry, and are bounded by $|H^{\prime}(\Delta)|=\sqrt{2}|K^{\prime}(\Delta)|$ since
$\int_{-\infty}^{H(\Delta)}p_{\Delta}\big(x_{1},H(\Delta)\big)\mathrm{d}x_{1}=\frac{e^{-H(\Delta)^{2}/2r_{n}}}{\sqrt{2\pi r_{n}}}\mathbb{P}\big(X_{1}\leq H(\Delta)\big|X_{2}=H(\Delta)\big)\leq 1$
uniformly over $\Delta$. As for the double integral in (4.2), differentiating the bivariate Gaussian density and using the triangle inequality gives
$\ll|K^{\prime}(\Delta)|\bigg(\frac{|K(\Delta)|}{(r_{n}^{2}-K(\Delta)^{2})}\mathbb{P}\big(X\leq H(\Delta)\big)+\frac{\mathbb{E}\big[|X_{1}X_{2}|K(\Delta)\mathbf{1}\big\{X\leq H(\Delta)\big\}\big]}{(r_{n}^{2}-K(\Delta)^{2})}$
(4.16) $+\frac{|K(\Delta)|}{(r_{n}^{2}-K(\Delta)^{2})^{2}}\mathbb{E}\big[\mathbf{1}\big\{X\leq H(\Delta)\big\}(K(\Delta)|X_{1}X_{2}|+r_{n}|X_{1}|+r_{n}|X_{2}|)\big]\bigg)$
where $X=\max(X_{1},X_{2})$. We prove that the right-hand side in (4.2) is bounded by a constant times $r_{n}^{2}|K^{\prime}(\Delta)|$ using rather crude bounds. To begin with, we can safely ignore the $(r_{n}^{2}-K(\Delta)^{2})^{-1}$ terms since
$r_{n}^{2}-K(\Delta)^{2}\geq r_{n}^{2}-(r_{n}-1)^{2}\geq 1,\quad\text{ if $|\Delta|>en^{-\delta}=e^{-r_{n}+1}$},\,r_{n}\geq 1.$
Secondly, we bound the probability term in (4.2) as well as the indicators in the expectations therein by one. Combined with the fact that $K(\Delta)\ll r_{n}$ in this range, this yields the simple bound
$\ll r_{n}^{2}|K^{\prime}(\Delta)|\cdot\Big(\mathbb{E}\big[|X_{1}X_{2}|\big]+\mathbb{E}\big[|X_{1}|\big]+\mathbb{E}\big[|X_{2}|\big]\Big).$
The remaining expectations can be estimated using the Cauchy-Schwarz inequality
$\mathbb{E}[|X_{1}X_{2}|]\ll r_{n},\quad\mathbb{E}[|X_{i}|]\ll\sqrt{r_{n}},\,i\in\{1,2\}.$
By combining the estimates above, we may conclude that
$\sup_{\Delta\in[\Delta_{n},1]}\partial_{\Delta}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\ll\sup_{\Delta\in[\Delta_{n},1]}\Big(r_{n}^{3}|K^{\prime}(\Delta)|+|K^{\prime}(\Delta)|\Big)\leq n^{\delta}r_{n}^{3}$
(where we used (4.13) to bound $|K^{\prime}(\Delta)|$). It follows that from (4.14),
(4.17) $\int_{\Delta_{n}}^{1}e^{in\Delta}e^{2K(\Delta)}\mathbb{Q}_{\Delta}(\mathcal{I}_{\leq r_{n}})\mathrm{d}\Delta\ll r_{n}^{3}n^{3\delta-1}=o\big((\log n)^{-2}\big),$
and in turn that
$e^{-2r_{n}}\iint_{(\mathrm{I})}e^{in(\theta_{2}-\theta_{1})}\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(\theta_{1})+X_{r_{n}}(\theta_{2}))}\mathcal{I}_{\leq r_{n}}\Big]\mathrm{d}\theta_{1}\mathrm{d}\theta_{2}=o\big((\log n)^{-2}\big).$
4.3. Dominant contribution
Having shown that $\mathbb{E}\mathcal{C}_{n,t}(\mathrm{I})$ can be discarded, it remains to bound the contribution coming from $\mathbb{E}\mathcal{C}_{n,t}(\mathrm{II})$. We prove the following bound.
Proposition 4.2.
Let $A>0$ and $\mathcal{E}_{n,t}^{(A)}$ be as in the statement of Lemma 4.2. Then
$\limsup_{t\to\infty}\mathbb{E}\big[\mathcal{C}_{n,t}\mathrm{(II)}\big]\ll_{A}\frac{1}{(\log n)^{2}}$
for large enough $n$.
We begin by introducing the auxiliary function
(4.18) $\mathcal{F}_{n,t}(\Delta)=te^{-2t}\cdot\mathbb{E}\Big[\mathbf{1}\big\{0,\Delta\in\mathcal{G}_{n,t}^{(A)}\big\}e^{\sqrt{2}(X_{t}(0)+X_{t}(\Delta))}\Big]$
noting that Fubini’s theorem and the triangle inequality yield
(4.19) $\mathbb{E}\big[\mathcal{C}_{n,t}(\mathrm{II})\big]\ll\int_{0}^{\Delta_{n}}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta.$
It therefore suffices to prove that the right-hand side is $\ll(\log n)^{-2}$ as $t\to\infty$. The bulk of the work lies in proving the following uniform upper bound for $\mathcal{F}_{n,t}$ whose proof is deferred to Section 4.3.
To state it, we introducing the branching time
${r_{\Delta}}=t\land\big(\log|\Delta|^{-1}\lor r_{n}\big),\quad\text{ for }\Delta\in[e^{-t},\Delta_{n}]$
This should essentially be viewed as $\log|\Delta|^{-1}$ on a first reading. The minimum and maximum taken in the definition only serve to make $r_{\Delta}$ well-behaved near the “boundary” (namely $\Delta\approx e^{-t}$ and $\Delta\approx\Delta_{n}$).
Proposition 4.3.
Let $\mathcal{F}_{n,t}$ be defined as in Equation (4.18). Then for large enough $t,n>0$, there exists a constant $C(A)=C>0$ such that $\mathcal{F}_{n,t}(\Delta)$ satisfies the bound
(4.20) $\mathcal{F}_{n,t}(\Delta)\leq C\frac{1}{r_{n}^{2}}\frac{t}{(t-{r_{\Delta}})+1}\bigg(\frac{\log(r_{\Delta}-r_{n})^{5}}{(r_{\Delta}-r_{n})^{2}}\land 1\bigg)\frac{1}{\Delta},\quad\text{ for }\Delta\in[e^{-t},\Delta_{n}]$
uniformly in $\Delta$ in said range, and
(4.21) $\mathcal{F}_{n,t}(\Delta)\leq Cf(n)\frac{e^{t}}{\sqrt{t}},\quad\text{ for }\Delta\in[0,e^{-t})$
for some positive function $f$ of $n$.
Proof of Proposition 4.2 assuming Proposition 4.3.
For large enough fixed $n>0$, we can use Proposition 4.3 to bound the integral in (4.19), beginning with the range $0,e^{-t}]$ where
[…]
On $e^{-t},\Delta_{n}]$, we use the bound in (4.20) to write
$\int_{e^{-t}}^{\Delta_{n}}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta\ll_{A}\frac{1}{{r_{n}}^{2}}\int_{e^{-t}}^{\Delta_{n}}\frac{t}{(t-{r_{\Delta}})+1}\bigg(\frac{\log(r_{\Delta}-r_{n})^{5}}{(r_{\Delta}-r_{n})^{2}}\land 1\bigg)\frac{1}{\Delta}\mathrm{d}\Delta.$
We first take care of the range where $r_{\Delta}$ is close to $r_{n}$. On $\Delta_{n}/e^{3},\Delta_{n}]$, say, we can bound the integral by
$\ll\frac{1}{r_{n}^{2}}\frac{t}{t-r_{n}}\int_{\Delta_{n}/e^{3}}^{\Delta_{n}}\frac{\mathrm{d}\Delta}{\Delta}\ll\frac{1}{r_{n}^{2}}\int_{r_{n}-1}^{r_{n}+2}\mathrm{d}\log(\Delta^{-1})\ll\frac{1}{r_{n}^{2}}.$
Using the substitution $u=r_{\Delta}$, we split the remaining range into $u\in[r_{n}+2,t/2)$ and $u\in[t/2,t)$. The contribution from the latter is
$\ll\frac{1}{r_{n}^{2}}\int_{t/2}^{t}\frac{t}{(t-r_{\Delta})+1}\frac{\log(u-r_{n})^{5}}{(u-r_{n})^{2}+1}\mathrm{d}u$
$\ll t\log(t)^{5}\int_{t/2}^{t-1}\frac{1}{u^{2}(t-u)}\mathrm{d}u$
$\ll\frac{\log(t)^{6}}{t}\to 0\text{ as }t\to\infty.$
On the former, $t/(t-u)\leq 2$ and we conclude that
$\limsup_{t\to\infty}\int_{e^{-t}}^{\Delta_{n}}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta$ $\ll_{A}\frac{1}{r_{n}^{2}}\int_{r_{n}+2}^{t/2}\frac{\log(u-r_{n})^{4}}{(u-r_{n})^{2}}\mathrm{d}u$
$\ll_{A}\frac{1}{r_{n}^{2}}\Big(1+\int_{1}^{\infty}\frac{\log(x)^{4}}{x^{2}}\mathrm{d}x\Big).$
Remark 4.4.
For any $a>b>\delta$, the same argument shows that
$\lim_{t\to\infty}\int_{n^{-a}}^{n^{-b}}e^{in\Delta}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta\ll\frac{1}{r_{n}^{2}}\int_{b\log n}^{a\log n}\frac{\log(u-r_{n})^{4}}{(u-r_{n}+1)^{2}}\mathrm{d}u\ll\frac{1}{(\log n)^{2}}$
whereas for any $s_{1}(n)\leq s_{2}(n)$ both tending to $\infty$ in $n$,
$\lim_{t\to\infty}\int_{n^{-s_{2}(n)}}^{n^{-s_{1}(n)}}e^{in\Delta}\mathcal{F}_{n,t}(\Delta)\mathrm{d}\Delta=o\bigg(\frac{1}{(\log n)^{2}}\bigg).$
This suggests that the main contribution to $\mathbb{E}[\mathcal{C}_{n,t}\mathrm{(II)}]$ (and in turn $\mathbb{E}[|c_{n,t}|^{2};\mathcal{E}_{n,t}^{(A)}]$) comes from pairs $(\theta_{1},\theta_{2})$ for which $\Delta\approx n^{-a}$ for all scales $a>\delta$.
5. Upper bound for $\mathcal{F}_{n,t}(\Delta)$
This section proves Proposition 4.3. To begin with, by independence of increments, the expectation in the definition of $\mathcal{F}_{n,t}$ splits into
(5.1) $\mathcal{F}_{n,t}(\Delta)=te^{-2t}\mathbb{E}\Big[e^{\sqrt{2}(X_{r_{n}}(0)+X_{r_{n}}(\Delta))}\mathbf{1}\big\{X_{r_{n}}(*)\leq m(r_{n})+A,*\in\{0,\Delta\}\big\}\Big]$
$\times\mathbb{E}\big[e^{\sqrt{2}(X_{t-r_{n}}^{(r_{n})}(0)+X_{t-r_{n}}^{(r_{n})}(\Delta))}\mathbf{1}\big\{X_{s}^{(r_{n})}(*)\leq U(s)+A,\forall s\in[0,t-r_{n}],*\in\{0,\Delta\}\big\}\Big].$
By Girsanov’s theorem, the first of the two expectations is bounded above by
$e^{2K_{r_{n}}(\Delta)+2r_{n}}\mathbb{Q}_{\Delta}\big(X_{r_{n}}(0)\leq m(r_{n})+A\big),$
where $\mathbb{Q}_{\Delta}$ is a probability measure under which the mean of $X_{r_{n}}(0)$ is shifted by $\sqrt{2}(r_{n}+K_{r_{n}}(\Delta))$ (see Equation 4.9). Since the latter is also Gaussian with variance $r_{n}$, it follows from that the above is
(5.2) $e^{2K_{r_{n}}(\Delta)+2r_{n}}\mathbb{P}\Big(\mathcal{N}\big(0,r_{n}\big)\leq-\sqrt{2}K_{r_{n}}(\Delta)-\frac{3}{2\sqrt{2}}\log(r_{n})+A\Big).$
For $\Delta\leq\Delta_{n}$, $K_{r_{n}}(\Delta)=r_{n}+O(1)$ (cf.  Equation (2.2)) and (5.2) is therefore bounded by
(5.3) $\ll e^{4r_{n}}\int_{-\infty}^{O_{A}(1)-\sqrt{2}r_{n}-\frac{3}{2\sqrt{2}}\log r_{n}}\frac{e^{-u^{2}/(2r_{n})}}{\sqrt{r_{n}}}\mathrm{d}u\ll_{A}\frac{e^{3r_{n}}}{r_{n}^{2}}$
for large enough $n\geq 0$.
We now move on to bounding
(5.4) $\mathbb{E}\big[e^{\sqrt{2}(X_{t-r_{n}}^{(r_{n})}(0)+X_{t-r_{n}}^{(r_{n})}(\Delta))}\mathbf{1}\big\{X_{s}^{(r_{n})}(*)\leq U(s)+A,\forall s\in[0,t-r_{n}],*\in\{0,\Delta\}\big\}\Big]$
beginning with another application of Girsanov’s theorem, by which the above equals
(5.5) $e^{2(t-r_{n})+2(K_{t}(\Delta)-K_{r_{n}}(\Delta))}$
$\times\mathbb{P}$ $\Big(X_{s}^{(r_{n})}(*)\leq A+{U}(s)-2\sqrt{2}s-\sqrt{2}\big(K_{s+r_{n}}(\Delta)-K_{r_{n}}(\Delta)\big),\begin{subarray}{c}\forall s\in[0,t-r_{n}],\\ *\in\{0,\Delta\}\end{subarray}\Big).$
The goal will be to prove that this quantity satisfies the bound
$\ll_{A}\frac{e^{2(t-r_{n})+(r_{\Delta}-r_{n})}\cdot\log(r_{\Delta}-r_{n})^{5}}{(r_{\Delta}-r_{n})^{2}\big((t-r_{\Delta})+1\big)},$
which, combined with (5.3) yields the desired estimate for $\mathcal{F}_{n,t}(\Delta)$, namely
$\mathcal{F}_{n,t}(\Delta)\ll_{A}\frac{1}{r_{n}^{2}}\frac{1}{\Delta}\cdot\frac{\log(r_{\Delta}-r_{n})^{5}}{(r_{\Delta}-r_{n})^{2}}\frac{t}{\big((t-r_{\Delta})+1\big)}.$
To estimate the probability in (5.5), we argue similarly to the proof of Proposition 2.6 in [23] but require a slightly sharper estimate, which we obtain by utilizing the lower order terms in $U(s)$. In particular, we will need the following “linearized” version of $U(s)$.
Recall that
$r_{\Delta}:=t\land\big((\log|\Delta|^{-1})\lor r_{n}\big),\quad\text{$r_{\Delta}\in[r_{n},t]$ for $\Delta\in[0,\Delta_{n}]$}.$
Whenever $r_{\Delta}-r_{n}\geq e$, we define the slope
$\alpha_{\Delta}=\alpha=\bigg(\sqrt{2}-\frac{\log(r_{\Delta}-r_{n})}{2\sqrt{2}(r_{\Delta}-r_{n})}\bigg),$
so that
(5.6) $U(s)\leq\alpha s+\frac{4}{\sqrt{2}}{\log\log(r_{\Delta}-r_{n})}\,\text{ for all $s\in[e,r_{\Delta}-r_{n}]$},$
with equality at $s=r_{\Delta}-r_{n}$.
We are now equipped to bound the probability in (5.5), assuming for now that $(r_{\Delta}-r_{n})\geq e$ so that $U(r_{\Delta}-r_{n})$ is well-defined. Observe first that by symmetry, we can add the restriction $X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)\leq X_{r_{\Delta}-r_{n}}^{(r_{n})}(\Delta)$ to the event in (5.5) since doing so can only change its probability by a factor of $1/2$. Secondly, we can use the estimate
$K_{s+r_{n}}(\Delta)-K_{r_{n}}(\Delta)=s\land(r_{\Delta}-r_{n})+O(1),$
in (5.5), which follows from Equation (2.5). It follows that there exists a constant $C(A)=C$ for which the intersection of $\{X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)\leq X_{r_{\Delta}-r_{n}}^{(r_{n})}(\Delta)\}$ with the events in (5.5) is contained in $E_{1}\cap E_{2}\cap E_{3}$, where
$E_{1}$ $:=\Big\{X_{s}^{(r_{n})}(0)\leq C+\overline{U}(s),\forall s\in[0,r_{\Delta}-r_{n}]\Big\}$
$E_{2}$ $:=\Big\{X_{s}^{(r_{\Delta})}(0)\leq C+\overline{U}(r_{\Delta}-r_{n})-X_{r_{\Delta}-r_{n}}^{(r_{n})}(0),\forall s\in[0,t-r_{\Delta}]\Big\}$
$E_{3}$ $:=\Big\{X_{s}^{(r_{\Delta})}(\Delta)\leq C+\overline{U}(r_{\Delta}-r_{n})-X_{r_{\Delta}-r_{n}}^{(r_{n})}(0),\forall s\in[0,t-r_{\Delta}]\Big\}$
and $\overline{U}(s):=U(s)-2\sqrt{2}s$.
Note that to get the inequality in $E_{3}$, we used (and subsequently dropped) the restriction $X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)\leq X_{r_{\Delta}-r_{n}}^{(r_{\Delta})}(\Delta)$. We have also used the fact that $\overline{U}$ is decreasing on $e,\infty)$.
By Equation (2.10) we have
$\mathbb{P}\Big(E_{2}\cap E_{3}\,\Big|\,X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)=z\Big)\ll\frac{(C+\overline{U}(r_{\Delta}-r_{n})-z+1)^{2}}{((t-r_{\Delta})+1)}.$
Furthermore, by (5.6),
$\mathbb{P}\Big(E_{1}\Big|X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)=z\Big)$
$\leq\mathbb{P}\bigg(X_{s}^{(r_{n})}(0)\leq C+\frac{4}{\sqrt{2}}{\log\log(r_{\Delta}-r_{n})}+(\alpha-2\sqrt{2})s,$
$\hskip 200.74998pt\forall s\in[e,r_{\Delta}-r_{n}]\Big|X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)=z\bigg),$
and we can estimate the right-hand side using Equation (2.11). Indeed, it is bounded by the probability that
$Z_{s}:=C+\frac{4}{\sqrt{2}}\log\log(r_{\Delta}-r_{n})+\bigg(\alpha-2\sqrt{2}-\frac{z}{(r_{\Delta}-r_{n})}\bigg)s-X_{s}^{(r_{n})}(0)+\frac{z}{(r_{\Delta}-r_{n})}s,$
which is a Brownian bridge from $C+(4/\sqrt{2})\log\log(r_{\Delta}-r_{n})$ to $(C+\overline{U}(r_{\Delta}-r_{n})-z)$, remains positive at all times $s\in[e,r_{\Delta}-r_{n}]$. This gives
$\ll\frac{(C+\log(r_{\Delta}-r_{n}))(C+\overline{U}(r_{\Delta}-r_{n})-z+1)}{((r_{\Delta}-r_{n})+1)},$
where we bounded the iterated logarithm by $\log(r_{\Delta}-r_{n})$ for simplicity.
By integrating over the value of $C+\overline{U}(r_{\Delta}-r_{n})-X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)$, we can then bound the quantity in (5.5) by
$e^{2(t-r_{n})+2(K_{t}(\Delta)-K_{r_{n}}(\Delta))}\mathbb{E}\bigg[\mathbf{1}_{E_{1}}\mathbb{P}\Big(E_{2}\cap E_{3}\,\Big|\,X_{r_{\Delta}-r_{n}}^{(r_{n})}(0)\Big)\bigg]$
$\ll e^{2(t-r_{n})+2(r_{\Delta}-r_{n})}$
$\times\int_{0}^{\infty}\frac{e^{-\frac{(x-C-\overline{U}(r_{\Delta}-r_{n}))^{2}}{2(r_{\Delta}-r_{n})}}}{\sqrt{r_{\Delta}-r_{n}}}\frac{(C+\log(r_{\Delta}-r_{n}))(x+1)(x+1)^{2}}{\big((r_{\Delta}-r_{n})+1\big)\big((t-r_{\Delta})+1\big)}\mathrm{d}x$
(5.7) $\ll_{A}e^{2(t-r_{n})+2(r_{\Delta}-r_{n})}\frac{\log(r_{\Delta}-r_{n})e^{-(\overline{U}(r_{\Delta}-r_{n})+C)^{2}/2(r_{\Delta}-r_{n})}}{(r_{\Delta}-r_{n})^{3/2}\big((t-r_{\Delta})+1\big)}.$
Recalling that
$\overline{U}(s)=-\sqrt{2}s-\frac{\log s}{2\sqrt{2}}+\frac{4}{\sqrt{2}}\log\log s,$
expanding the square in the exponential in Equation (5) gives
$\ll_{A}\frac{e^{2(t-r_{n})+(r_{\Delta}-r_{n})}\cdot\log(r_{\Delta}-r_{n})^{5}}{(r_{\Delta}-r_{n})^{2}\big((t-r_{\Delta})+1\big)},$
If $r_{\Delta}-r_{n}<e$, repeating the argument while omitting $E_{1}$ yields the desired bound, noting that $e^{r_{\Delta}}\asymp e^{r_{n}}$ in this range. Lastly, the bound in (4.21) follows immediately from the above by noting that $r_{\Delta}=t$ when $\Delta\leq e^{-t}$, by definition.
Funding
L.-P. A.’s research is supported in part by the NSF grant DMS-2153803 and the EPSRC grant EP/Z535990/1. J.H. supported by the EPSRC Centre for Doctoral Training in Mathematics of Random Systems: Analysis, Modelling and Simulation (EP/S023925/1). For the purpose of open access, the authors have applied a CC BY public copyright licence to any author accepted manuscript arising from this submission.